OpenStudy (anonymous):

the funtion g is defined by g(x)=(7+3x)/(6+x) find g(n+1)

OpenStudy (ivettef365):

substitute the x for (n+1) on the function g(x)

OpenStudy (anonymous):

g(n+1)=(7+3(n+1))/(6+(n+1))

OpenStudy (anonymous):

what do you do after you substitute

OpenStudy (ivettef365):

solve for n

OpenStudy (ivettef365):

7 + 3n + 3 3n + 10 ---------- = ------------- 6 + n + 1 n + 7

OpenStudy (anonymous):

ya thats what i got would that be the answer then

OpenStudy (ivettef365):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

one more question if the answer came out to something like 4n/2n could you simplify it or do you have to leave it